OpenStudy (anonymous):

show that d/dx ((sin^2x+1)/(cot(x))-(cos^2(x)+1)/tan(x)) = 2sec^2(x)csc^2(x)

OpenStudy (anonymous):

\[\frac{d}{dx}\left[\frac{\sin^2x+1}{\cot x}-\frac{\cos^2x+1}{\tan x}\right]=2\sec^2x\csc^2x~~?\]